OpenStudy (anonymous):

If tan A = m/m-1 , tan B = 1/2m-1 . St (A-B) = π /4.

OpenStudy (anonymous):

St ??

OpenStudy (anonymous):

show or state that !

hartnn (hartnn):

try this: (tan A-tan B)/(1+tanAtanB).....find this value by substituting from above equation...if u get answer as 1,u can use tan (A-B)=(tan A-tan B)/(1+tanAtanB)=1= tan pi/4...so A-B=pi/4

OpenStudy (anonymous):

Find tan(A - B)..

OpenStudy (anonymous):

okay , let me try :)

OpenStudy (anonymous):

\[\tan(a - B ) = \frac{\tan(A)- \tan(B)}{1 + \tan(A) \tan(B)}\]

OpenStudy (anonymous):